obsidian1:

Which equivalent expression will be generated by applying the Distributive Property and combining like terms in the expression 11 + 4(x + 2y + 4)? A. 4x + 2y + 27 B. 4x + 8y + 11 C. 4x + 8y + 27 D. 4x + 8y + 10

prettygirlgang:

wait 1 sec

supie:

\(11+4(x+2y+4)\) So first you distribute `4` to everything inside the parenthesis So \(11+(4)(x)+(4)(2y)+(4)(4)\) which equals \(11+4x+8y+16\) So then you combine the like terms which are `11 and 16`, So re-write the equation, \((4x)+(8y)+(11+16)\) then add 11 and 16, remove the parenthesis then you'll have your answer

supie:

@obsidian1,,,

obsidian1:

ok thats confusing

supie:

\(11+4(x+2y+4)\) So you have to distribute the `4` to everything inside the parenthesis which is `x` `2y` and `4` So \(4*x=4x\\4*2y=8y\\4*4=12\) So \(11+4(x+2y+4)\longrightarrow 11+(4)(x)+(4)(2y)+(4)(4)\) So we already solved those \(11+(4)(x)+(4)(2y)+(4)(4)\longrightarrow 11+4x+8y+16\) So basically, what we just did was distribute the '4' to everything inside the parenthesis so it became \(11+\color{red}{(4)(x)}+\color{red}{(4)(2y)}+\color{red}{(4)(4)}\) Which is; \(4*x=4x\\4*2y=8y\\4*4=12\) so we are solving everything that was red So then we got \(11+4x+8y+16\) So now we combine the like terms inside the equation \(11+4x+8y+16\) So basically, since there is no like terms together, we cant combine anything So we have to move the numbers so that we can combine the like terms So we can do \((4x)+(8y)+(11+16)\) (We did that so we can get `11` and `16` next to each other because those are the only like terms, so we had to get them next to each other so that we can add them together - so we basically just moved the numbers...if that makes sense) So then we combine the like terms which is `11` and `16` and we get ___? So basically, all you have to do is solve that 11+16, then remove the parenthesis from \(\color{red}{(}4x\color{red}{)}+\color{red}{(}8y\color{red}{)}+\color{red}{(}11+16\color{red}{)}\) (fisrt you have to solve 11+16) then you'll have your answer. of any of that doesn't make sense lmk I tried to make it make the most sense as I could :/ (So the answer would be 4x+8y+?( `?` stands for what `11+16` equals))